OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#5796 — Optical loop (LIB59 SEC59 FAC59 BOI59 BLE59 PIL59)
Scheduled Maintenance Report for Web Cloud
Completed
We are doing working tonight at midnight in order to replace the optical fibers. An interruption of a few minutes is expected on each of the sites.
We will update regularly this task in order to keep you informed of the task's evolution.

Update(s):

Date: 2011-09-20 00:14:51 UTC
The site of BOI59 is UP. We will finish with the site BLE59.

Date: 2011-09-20 00:14:06 UTC
The site of FAC59 is UP.We will continue with the site of BOI59 in a few minutes.

Date: 2011-09-20 00:13:09 UTC
The site of SEC59 is UP.We will continue with the site of FAC59 in a moment.

Date: 2011-09-20 00:12:00 UTC
The site of LIB59 is UP.We will continue with the site of SEC59 in a few minutes.

Date: 2011-09-19 22:51:48 UTC
The site of LIB59 will be disconnected in a few minutes.

Date: 2011-09-19 22:50:57 UTC
The PIL site is UP. We are moving forward with works!

Date: 2011-09-19 22:49:56 UTC
PIL59 fiber is reconnected, we reconfigure the routing.

Date: 2011-09-19 22:49:21 UTC
We are unplugging PIL59 in order to rewire it. A few minutes cut is expected.
Posted Sep 19, 2011 - 19:00 UTC